# Efficient Shapley Value-based Non-Uniform Pruning of Large Language Models

### v2

@@ Abstract
Pruning large language models (LLMs) is a promising solution for reducing model sizes and computational complexity while preserving performance. Traditional layer-wise pruning methods often adopt a uniform sparsity approach across all layers, which leads to suboptimal performance due to the varying significance of individual transformer layers within the model not being accounted for. To this end, we propose the Shapley Value-based Non-Uniform Pruning (SV-NUP) method for LLMs. This approach quantifies the contribution of each transformer layer to the overall model performance, enabling the assignment of tailored pruning budgets to different layers to retain critical parameters. To further improve efficiency, we design the Sliding Window-based Shapley Value approximation method. It substantially reduces computational overhead compared to exact SV calculation methods. Extensive experiments on various LLMs including LLaMA-v1, LLaMA-v2 and OPT demonstrate the effectiveness of the proposed approach. The results reveal that non-uniform pruning significantly enhances the performance of pruned models. Notably, SV-NUP achieves a reduction in perplexity (PPL) of 18.01% and 19.55% on LLaMA-7B and LLaMA-13B, respectively, compared to SparseGPT at 70% sparsity.
@@ 1 Introduction
Large language models (LLMs) have emerged as a transformative technology, demonstrating remarkable capabilities in natural language understanding and generation tasks Brown et al. (2020). LLMs have showcased significant adaptability through fine-tuning, enabling their deployment in highly specialized applications. These advantages underscore the critical role of LLMs in solving real-world challenges Patterson et al. (2021). Despite these capabilities, the deployment of LLMs has been hindered by their immense computational demands. Modern LLMs often consist of billions or even trillions of parameters Luccioni et al. (2023), as seen in models like GPT-3 (175 billion parameters) and PaLM (540 billion parameters). The immense scale incurs significant memory, storage and power costs, making it challenging to run these models on resource-constrained devices. To address these issues, researchers have increasingly turned to model compression techniques, particularly pruning, to reduce the size and computational requirements of LLMs, while retaining their performance Frantar and Alistarh (2023).
@@
Pruning LeCun et al. (1989); Han et al. (2015b, a) is one of the most prominent model compression techniques. It aims to remove redundant or less important parameters from neural networks, thereby reducing their sizes and computational complexity. Recent advances in pruning have progressed from unstructured sparsity (i.e., individual weight removal) to structured sparsity (i.e., eliminating entire neurons, heads or layers), thereby enabling hardware-efficient implementation Frantar and Alistarh (2023); Jaiswal et al. (2024a); Ma et al. (2023); Xia et al. (2023). For LLMs, several state-of-the-art methods have been proposed, including magnitude pruning, lottery ticket hypothesis approaches, and structured pruning based on attention mechanisms. These methods demonstrate the potential of pruning to enable LLM deployment on resource-constrained devices without causing substantial performance degradation.
@@
Existing pruning methods mostly adopt a layer-wise approach to deal with large-scale parameters in LLMs. Layer-wise pruning is a uniform sparsity strategy, applying the same pruning ratio across all layers of an LLM. While simple to implement, this approach overlooks the inherent differences in the contributions of different layers to the overall performance of the model. Thus, it can only find the local optimal pruning solution, but not the global optimal solution. Empirical evidence also suggests that certain layers are more critical than others, and uniformly pruning across all layers may lead to the removal of essential parameters, ultimately impairing the pruned model’s performance Mocanu et al. (2018); Yin et al. (2023).
@@
To address this limitation, we leverage the concept of Shapley value from cooperative game theory and propose a novel Shapley Value-based Non-Uniform Pruning (SV-NUP) method for LLMs, which was originally designed to fairly assess the contributions of multiple players in a game. In this study, we treat each Transformer in an LLM as a “player”. Based on this idea, we evaluate the contribution of each Transformer to the overall performance of an LLM. The results are further used as a basis to allocate customized sparsity ratios to each Transformer, prioritizing the preservation of parameters in more important ones. This approach not only improves the performance of the pruned LLMs, but also introduces a theoretical foundation for sparsity allocation, moving beyond heuristic methods. We further propose a Sliding Window-based Shapley Value approximation method (SWSV) to reduce the computational cost.
@@
We conduct extensive experiments to evaluate the performance of SV-NUP for pruning LLMs across a spectrum of LLMs, including LLaMA-v1, LLaMA-v2 and OPT. SV-NUP achieves a reduction in perplexity (PPL) of 18.01% and 19.55% on LLaMA-7B and LLaMA-13B, respectively, compared to SparseGPT at 70% sparsity. The results validate our hypothesis that non-uniform sparsity allocation can significantly enhance the performance of pruned models, paving the way for more efficient and practical LLM deployment on resource-constrained devices.
@@ 2 Related Work
Uniform Pruning. Traditional pruning requires a round of re-training to restore performance, which poses significant challenges for LLMs. Researchers have developed pruning algorithms specifically tailored for LLM compression. For instance, Ma et al. (2023) investigated structured sparse LLMs by applying Taylor pruning to remove entire weight rows, followed by LoRA fine-tuning Hu et al. (2021). In recent years, the focus has shifted toward unstructured pruning which eliminates the need for fine-tuning. SparseGPT Frantar and Alistarh (2023) employs the Hessian inverse for pruning, followed by weight updates to reduce reconstruction errors between dense and sparse weights. Wanda Sun et al. (2023) introduced a criterion that incorporates weight magnitude and input activations to preserve outlier features.
Non-uniform Pruning. Uniform layerwise sparsity is commonly used for pruning language models Zhu and Gupta (2017); Gale et al. (2019), with several studies demonstrating its effectiveness in LLM pruning Sanh et al. (2020); Kurtic et al. (2022). However, there is a growing body of work exploring non-uniform layerwise sparsity, primarily in the context of vision models. For example, Mocanu et al. (2016) proposed a non-uniform, scale-free topology inspired by graph theory, which outperforms dense counterparts when applied to restricted Boltzmann machines. Subsequent work has improved the scalability of this approach by leveraging Erdős–Rényi graphs ERDdS and R&wi (1959), extending the method to fully connected layers Mocanu et al. (2018) and convolutional layers Evci et al. (2020); Liu et al. (2022) to achieve data-free and feedforward-free layerwise sparsity. Another approach to non-uniform sparsity involves applying a global threshold across all layers Frankle and Carbin (2018); Lee et al. (2018); Wang et al. (2020); Liu et al. (2021). However, global pruning has been found to be computationally expensive and ineffective when applied to LLMs.
Analyzing LLMs. The authors in Shim et al. (2022) analyzed the contributions of various components in LLMs and their impact on overall performance. Gromov et al. (2024) explored the role of deep layers in LLMs through layer pruning, providing insights into how different layers in relation to model performance. Michel et al. (2019) examined the redundancy of attention heads in transformer-based models, demonstrating that many attention heads can be pruned without significant performance degradation. Clark (2019) investigated the behavior of individual attention heads in BERT, revealing that each head serves a distinct role in capturing different linguistic features. Probing techniques are widely used to analyze the internal representations of LLMs. For instance, Tenney (2019) employed probing tasks to examine the linguistic information captured by BERT, finding that different layers encode distinct types of linguistic features. Furthermore, Tenney et al. (2019) introduced a suite of probes to analyze the representations learned by contextualized word embeddings, offering insights into how syntactic and semantic information is distributed across layers. Recently, Zhang et al. (2024) used Shapley values to evaluate the importance of layers in LLMs, providing a faithful assessment of their contributions. However, these studies have not explored Transformer SV-based non-uniform pruning of LLMs. SV-NUP bridges this gap.
@@ 3 Motivation
In this section, we conduct empirical studies on three LLMs to demonstrate our motivation for allocating different pruning ratios for different Transformers. We select 3 comment LLMs: 1) TinyLLama, 2) LLama-7B and 3) Mistral-7B. We compare the difference between Transformers by mean and standard deviation (std) of the Magnitude method Lee et al. (2020), which is widely used in neural network pruning. Based on the results shown in Figure 1, we can make the following two observations.
@@ 3.1 Observations
(a) TinyLLaMA
(b) TinyLLaMA
(c) LLaMA-7B
(d) LLaMA-7B
(e) Mistral-7B
(f) Mistral-7B
Figure 1: Comparison of the mean and standard deviation (std) across different Transformers under Magnitude for 3 LLMs (TinyLLama, LLama-7B, and Mistral-7B). TinyLLama consists of 22 Transformers and the other 2 LLMs consist of the same 32 Transformers, where each Transformer consists of 7 innerlayers.
Observation 1
Different Transformers in LLMs exhibit significant variations in the mean and std of magnitude.
As the Transformer index increases, the mean magnitude decreases across all LLMs (Figures 1(a), 1(c) and 1(e)). Notably, the changes in LLaMA-7B and Mistral-7B are more pronounced compared to TinyLLaMa. The earlier (shallower) Transformers tend to have higher mean magnitudes. From the std of magnitude in Figures 1(b), 1(d) and 1(f), it can be observed that weights of shallow Transformers change dramatically. It is evident that the weights of shallow Transformers vary more significantly. Higher mean and std of magnitude values indicate that the correspondong Transformers have greater impact on LLM performance. These results underscore that different layers contribute unevenly to the overall model performance, suggesting that some Transformers are more sensitive to changes in parameter magnitudes.
Observation 2
The attention module of Transformers has a higher mean and std of magnitude than the feed-forward network (FFN) module.
Each Transformer consists of 7 inner layers. The first 4 correspond to the attention module, while the last 3 form the FFN module. In all the figures, the mean and std of magnitude in the attention module are higher than those in the FFN module. This finding indicates that the attention module is more sensitive compared to the FFN module to changes in parameter magnitude. For instance, in Figures 1(a) and 1(b), TinyLLaMa’s inner layer 4 shows a distinct dividing line: layers 1–4 form peaks, whereas layers 5–7 form troughs. The other three LLMs follow similar trends. As a result, the attention module warrants more attention during pruning.
@@ 3.2 Analysis and Implications
Conventional LLM pruning approaches that apply a uniform pruning ratio across all Transformers tend to overlook the aforementioned inherent diversity. For example, in Figure 1(c), the first few Transformers exhibit lower mean magnitudes, suggesting that they can withstand more aggressive pruning; whereas the latter Transformers show higher mean magnitudes and std, indicating that they should be pruned more conservatively to maintain performance. Applying the same pruning ratio to Transformers with high mean magnitudes or high std as those with lower values might remove critical capabilities from important Transformers, while preserving too many parameters in less significant ones.
This analysis clearly demonstrates the advantage of allocating different pruning ratios to different Transformers. By tailoring the pruning ratios based on the unique characteristics of each Transformer, pruning strategies can be optimized to retain crucial information in the model. Such a non-uniform pruning approach enables more efficient compression where less critical Transformers undergo more aggressive pruning, while vital Transformers are preserved, thereby enhancing the performance of the pruned model.
@@ 4 The Proposed SV-NUP Method
Motivated by these observations, we investigate the issue of adaptive sparsity for LLM pruning. The Shapley value, a concept derived from cooperative game theory, is widely used to fairly allocate contributions among multiple players in a game. This paper leverages the Shapley value to allocate pruning ratios across different Transformers. By treating each Transformer as a “player” in the model, we can compute the contribution of each Transformer to the overall performance of the LLM. This approach will allow us to assign pruning ratios dynamically, ensuring that more important Transformers (i.e., those with higher Shapley values) are pruned less aggressively, while less critical Transformers receive higher pruning ratios. We propose the Shapley Value-based Non-Uniform Pruning (SV-NUP) method to address this problem, which provides a theoretically grounded and equitable solution to the problem of non-uniform pruning. This method optimizes the trade-off between model size and performance, ensuring that critical Transformers are preserved, while less significant ones are pruned more aggressively.
@@
We consider post-training pruning of LLMs from a well-optimized model with weights $\mathbf{W}^{*}$ to a sparse version $\mathbf{W}$ with many 0 values under a given pruning ratio $\rho$. As LLMs contain hundreds of millions of parameters, we usually adopt layer-wise pruning. The neural network structure of LLMs is generally composed of many Transformers. We denote Transformers as the set $\mathcal{T}=\{1,2,...,T\}$. Conventional LLM pruning methods apply a uniform pruning ratio $\rho$ to all Transformers. According to the analysis in Section 3, we should allocate different pruning ratios $\rho_{t}$ to different Transformer $t\in\mathcal{T}$. The key to allocating pruning ratios is to calculate the contribution of Transformers of LLMs precisely. We use SV to analyze each Transformer contribution.
4.1 Transformer Contribution by Shapley Value
SV is a unique contribution allocation scheme that satisfies a set of fairness axioms. Here, we leverage the concept to calculate the weighted average of all marginal contributions of each Transformer in an LLM. To illustrate the calculation process, we consider a simple LLM with 3 Transformers as shown in Figure 2. Since all Transformers have the same structure, we can easily mask a Transformer during inference and do not affect SV calculation, as shown in the bottom right corner of Figure 2. We use the accuracy of an LLM as the basis for calculating the SVs.
@@ 4.1 Transformer Contribution by Shapley Value
Figure 2: Contribution analysis of 3 Transformers in an LLM. For intuitive presentation, we use accuracy to represent the contribution of different Transformers.
@@
We first identify each Transformer’s contribution when they participate individually, when 2 participate together, and when all 3 participate together. Particularly, as an LLM without any Transformer cannot function normally, we assume the accuracy to be 0. Then, we consider all possible combinations of Transformers and calculate their marginal values (e.g. what value does each Transformer add when Transformer 1 enters the LLM first, followed by Transformer 2, and then Transformer 3). Finally, we need to add them up and work out the SV (i.e. the average) for each Transformer.
@@
To formalize this process, an LLM consists of $\mathcal{T}$ Transformers and a value function $\nu$. Perplexity (PPL) is usually used to evaluate LLMs in LLM pruning JAISWAL et al. (2024b); Xia et al. (2024); Ma et al. (2023). The lower the value, the better the performance. Therefore, the value function is denoted as $\nu(\mathcal{T})=1/\text{PPL}(\mathcal{T})$. Particularly, if there is no Transformer, the value function must be 0, i.e., $\nu(\emptyset)=0$. Let $\mathcal{S}\subseteq\mathcal{T}$ denote a subset of $\mathcal{T}$. According to Shapley (1953), the SV of any Transformer $t\in\mathcal{T}$ is:
@@
where $w_{\mathcal{S}}=T\cdot\binom{T-1}{\mathcal{S}}=\frac{T!}{|\mathcal{S}|!(T-|\mathcal{S}|-1)!}$ is a coefficient. $\nu(\mathcal{S}\cup\{t\})-\nu(\mathcal{S})$ is known as Transformer $t$’s marginal contribution. Eq. (1) yields the following desirable properties:
  • Efficiency: The sum of all Shapley values equals to the total contribution towards an LLM by all Transformers, i.e., $\sum_{t\in\mathcal{T}}\phi_{t}=v(\mathcal{T})$. This ensures that all contributions from Transformers are fully allocated with no “leftover” or “missing” gain.
  • Symmetry: If two Transformers contribute equally to an LLM (i.e., they have the same marginal contribution in all subsets), they should have the same Shapley value, i.e., $v(\mathcal{S}\cup\{t\})=v(\mathcal{S}\cup\{i\}),\forall\mathcal{S}\subseteq\mathcal{T}\setminus\{t,i\}\Rightarrow\phi_{t}=\phi_{i}$. This ensures fairness, as Transformers with the same effect on an LLM are not favored over one another.
  • Dummy: If a Transformer does not add any marginal contribution to any subset of Transformers, its Shapley value is 0, i.e., $v(\mathcal{S}\cup\{t\})=v(\mathcal{S}),\forall\mathcal{S}\subseteq\mathcal{T}\Rightarrow\phi_{t}=0$. This ensures that Transformers that do not affect an LLM’s output receive no contribution value.
Eq. (1) involves calculating the marginal contribution for every subset $\mathcal{S}$, which results in a combinatorial explosion as the number of Transformers increases. Specifically, for $T$ Transformers, the total number of possible subsets $\mathcal{S}$ is $2^{T}$. For instance, with 32 Transformers in LLaMa-7B, the number of possible subsets becomes $2962^{32}=4,294,967,$ 296. The exact computation of Shapley values requires evaluating every subset that excludes the given Transformer, meaning that $32\times 2^{32}$ evaluations are needed. Each additional feature doubles the number of subsets that must be considered, causing an exponential increase in the number of computations. Consequently, this problem is NP-hard.

### v3

@@ Abstract
Pruning large language models (LLMs) is a promising solution for reducing model sizes and computational complexity while preserving performance. Traditional layer-wise pruning methods often adopt uniform sparsity at all Transformer layers, which leads to suboptimal performance due to the varying significance of transformer layers not being accounted for. To this end, we propose the Shapley Value-based Non-Uniform Pruning (SV-NUP) method for LLMs. This method quantifies the contribution of each transformer layer to the overall model performance, enabling the assignment of tailored pruning budgets to different layers to retain critical parameters. To further improve efficiency, we design the Sliding Window-based Shapley Value (SWSV) approximation method. It substantially reduces computational overhead compared to exact SV calculation methods. Extensive experiments on various LLMs including LLaMA-v1/v2/v3, and OPT demonstrate the effectiveness of SV-NUP. The results reveal that non-uniform pruning significantly enhances the performance of pruned models. Notably, SV-NUP achieves a reduction in perplexity (PPL) of 18.01% and 19.55% on LLaMA-7B and LLaMA-13B, respectively, compared to SparseGPT at 70% sparsity.
@@ 1 Introduction
Large language models (LLMs) have emerged as a transformative technology, demonstrating remarkable capabilities in natural language understanding and generation tasks [1]. LLMs have showcased significant adaptability through fine-tuning, enabling their deployment in highly specialized applications. These advantages underscore the critical role of LLMs in solving real-world challenges [2]. Despite these capabilities, the deployment of LLMs has been hindered by their immense computational demands. Modern LLMs often consist of billions or even trillions of parameters [3], as seen in models like GPT-3 (175 billion parameters) and PaLM (540 billion parameters). The immense scale incurs significant memory, storage and power costs, making it challenging to run these models on resource-constrained devices. To address these issues, researchers have increasingly turned to model compression techniques, particularly pruning, to reduce the size and computational requirements of LLMs, while retaining their performance [4].
@@
Pruning [5; 6; 7] is one of the most prominent model compression techniques. It aims to remove redundant or less important parameters from neural networks, thereby reducing their sizes and computational complexity. Recent advances in pruning have progressed from unstructured sparsity (i.e., individual weight removal) to structured sparsity (i.e., eliminating entire neurons, heads or layers), thereby enabling hardware-efficient implementation [4; 8; 9; 10]. For LLMs, several state-of-the-art methods have been proposed, including magnitude pruning, lottery ticket hypothesis approaches, and structured pruning based on attention mechanisms. These methods demonstrate the potential of pruning to enable LLM deployment on resource-constrained devices without causing substantial performance degradation.
Figure 1: Overview and effectiveness of SV-NUP: (a) Distribution of weight magnitude across Transformer layers; (b) Conceptual comparison of uniform pruning vs. SV-NUP; (c) PPL↓  (the smaller the better) comparison under different pruning strategies. The red number is the improvement percentage. Unlike uniform pruning, which ignores layer-wise importance and may degrade LLM performance, SV-NUP leverages Shapley value to estimate the contribution of each Transformer layer and allocates pruning ratios accordingly to better preserve model performance.
@@
Existing pruning methods often use a layer-wise strategy that applies uniform sparsity across all layers, ignoring their varying importance. While simple to implement, this approach overlooks the inherent differences in the contributions of different layers to the overall performance of the model. Thus, it can only find the local optimal pruning solution, but not the global optimal solution. Empirical evidence also suggests that certain layers are more critical than others, and uniformly pruning across all layers may lead to the removal of essential parameters, ultimately impairing the pruned model’s performance [11; 12].
@@
To address this limitation, we leverage the concept of Shapley value from cooperative game theory and propose a novel Shapley Value-based Non-Uniform Pruning (SV-NUP) method for LLMs, which was originally designed to fairly assess the contributions of multiple players in a game. In this study, we treat each Transformer layer in an LLM as a “player". Based on this idea, we evaluate the contribution of each Transformer layer to the overall performance of an LLM. The results are further used as a basis to allocate customized sparsity ratios to each Transformer layer, prioritizing the preservation of parameters in more important ones. This approach not only improves the performance of the pruned LLMs, but also introduces a theoretical foundation for sparsity allocation, moving beyond heuristic methods. We further propose a Sliding Window-based Shapley Value approximation method (SWSV) to reduce the computational cost. The overview of SV-NUP is shown in Figure 1.
@@
We conduct extensive experiments to evaluate the performance of SV-NUP for pruning LLMs across a spectrum of LLMs, including LLaMA-v1/v2/v3, and OPT. SV-NUP achieves a reduction in perplexity (PPL) of 18.01% and 19.55% on LLaMA-7B and LLaMA-13B, respectively, compared to SparseGPT at 70% sparsity. The results validate our hypothesis that non-uniform sparsity allocation can significantly enhance the performance of pruned models, paving the way for more efficient and practical LLM deployment on resource-constrained devices.
@@ 2 Related Work
LLM pruning often requires retraining to recover performance, which presents substantial challenges. While existing LLM-specific pruning methods primarily adopt uniform pruning strategies [9; 13; 4; 14], non-uniform layerwise sparsity has been extensively studied in vision models [15; 16; 11; 17; 18; 19; 20; 21; 22]. However, global pruning is often inefficient and computationally expensive for LLMs. Although prior work has analyzed the importance of LLM components [23; 24; 25; 26; 27; 28; 29], none have explored Shapley value (SV)-based non-uniform pruning. SV-NUP fills this gap. For a more detailed literature review, refer to Appendix A.
@@ 3 Motivation
In this section, we conduct empirical studies on three LLMs (TinyLLama, LLaMA-7B, and Mistral-7B) to motivate the need for allocating different pruning ratios across Transformer layers. We analyze the differences using the Magnitude method [30] and summarize each Transformer layer’s structure with the mean and standard deviation of innerlayer metrics. As shown in Figure 2, two key observations can be drawn from the results.
(a) TinyLLaMA with Mean
(b) LLaMA-7B with Mean
(c) Mistral-7B with Mean
(d) TinyLLaMA with Std
(e) LLaMA-7B with Std
(f) Mistral-7B with Std
Figure 2: Comparison of the mean and standard deviation (Std) across different Transformer layers under Magnitude. TinyLLaMA consists of 22 Transformer layers, LLama-7B and Mistral-7B consist of the same 32 Transformer layers, where each Transformer layer consists of 7 inner-layers.
@@ 3.1 Observations
Obervation 1 Different Transformer layers in LLMs exhibit significant variations in the mean and std of magnitude.
As the Transformer index increases, the mean magnitude decreases across all LLMs (Figures 2(a), 2(b) and 2(c)). Notably, the changes in LLaMA-7B and Mistral-7B are more pronounced compared to TinyLLaMa. The earlier (shallower) Transformer layers tend to have higher mean magnitudes. From the std of magnitude in Figures 2(d), 2(e) and 2(f), it can be observed that weights of shallow Transformer layers change dramatically. It is evident that the weights of shallow Transformer layers vary more significantly. Higher mean and std of magnitude values indicate that the corresponding Transformer layers have a greater impact on LLM performance. These results underscore that different layers contribute unevenly to the overall model performance, suggesting that some Transformer layers are more sensitive to changes in parameter magnitudes.
Obervation 2 The attention module of Transformer layers has a higher mean and std of magnitude than the feed-forward network (FFN) module.
Each Transformer layer consists of 7 inner layers. The first 4 correspond to the attention module, while the last 3 form the FFN module. In all the figures, the mean and std of magnitude in the attention module are higher than those in the FFN module. This finding indicates that the attention module is more sensitive compared to the FFN module to changes in parameter magnitude. For instance, in Figures 2(a) and 2(d), TinyLLaMa’s inner layer 4 shows a distinct dividing line: layers 1–4 form peaks, whereas layers 5–7 form troughs. The other three LLMs follow similar trends. As a result, the attention module warrants more attention during pruning.
@@ 3.2 Analysis and Implications
Conventional LLM pruning approaches apply uniform pruning across all Transformer layers, ignoring differences across Transformer layers. As shown in Figure 2(b), early Transformer layers exhibit lower mean magnitudes, suggesting that they can withstand more aggressive pruning; whereas the latter Transformer layers show higher mean magnitudes and std, indicating that they should be pruned more conservatively to maintain performance. Applying the same pruning ratio to Transformer layers with high mean magnitudes or high std as those with lower values might remove critical capabilities from important Transformer layers, while preserving too many parameters in less significant ones.
This analysis clearly demonstrates the advantage of allocating different pruning ratios to different Transformer layers. By tailoring the pruning ratios based on the unique characteristics of each Transformer layer, pruning strategies can be optimized to retain crucial information in the model. Such a non-uniform pruning approach enables more efficient compression where less critical Transformer layers undergo more aggressive pruning, while vital Transformer layers are preserved, thereby enhancing the performance of the pruned model.
@@ 4 The Proposed SV-NUP Method
Motivated by these observations, we investigate the issue of adaptive sparsity for LLM pruning. The Shapley value, a concept derived from cooperative game theory, is widely used to fairly allocate contributions among multiple players in a game. This paper leverages the Shapley value to allocate pruning ratios across different Transformer layers. By treating each Transformer layer as a “player" in LLMs, we can compute the contribution of each Transformer layer to the overall performance of an LLM. This approach will allow us to assign pruning ratios dynamically, ensuring that more important Transformer layers (i.e., those with higher Shapley values) are pruned less aggressively, while less critical Transformer layers receive higher pruning ratios. We propose SV-NUP to address this problem, which provides a theoretically grounded and equitable solution to non-uniform pruning. This method optimizes the trade-off between model size and performance, ensuring that critical Transformer layers are preserved, while less significant ones are pruned more aggressively.
@@
We consider post-training pruning of LLMs from a well-optimized model with weights $\mathbf{W}^{*}$ to a sparse version $\mathbf{W}$ with many 0 under a given pruning ratio $\rho$. As LLMs contain hundreds of millions of parameters, we usually adopt layer-wise pruning. The neural network structure of LLMs is generally composed of many Transformer layers. We denote Transformer layers as the set $\mathcal{T}=\{1,2,...,T\}$. Conventional LLM pruning methods apply a uniform pruning ratio $\rho$ to all Transformer layers. According to the analysis in Section 3, we should allocate different pruning ratios $\rho_{t}$ to different Transformer layer $t\in\mathcal{T}$. The key to allocating pruning ratios is to calculate the contribution of Transformer layers of LLMs precisely. We use SV to analyze each Transformer layer contribution.
4.1 Transformer Layer Contribution by Shapley Value
@@
Figure 3: Contribution analysis of 3 Transformer layers in an LLM. For intuitive presentation, we use accuracy to represent the contribution of different Transformer layers.
SV is a unique contribution allocation scheme that satisfies a set of fairness axioms. Here, we leverage the concept to calculate the weighted average of all marginal contributions of each Transformer layer in an LLM. To illustrate the calculation process, we consider a simple LLM with 3 Transformer layers as shown in Figure 3. Since all Transformer layers have the same structure, we can easily mask a Transformer layer during inference and do not affect SV calculation, as shown in the bottom right corner of Figure 3. We use the accuracy of an LLM as the basis for calculating the SVs.
@@
We first identify each Transformer layer’s contribution when they participate individually, when 2 participate together, and when all 3 participate together. Particularly, as an LLM without any Transformer layer cannot function normally, we assume the accuracy to be 0. Then, we consider all possible combinations of Transformer layers and calculate their marginal values (e.g. what value does each Transformer layer add when Transformer layer 1 enters the LLM first, followed by Transformer layer 2, and then Transformer layer 3). Finally, we need to add them up and work out the SV (i.e., the average) for each Transformer layer.
@@
To formalize this process, an LLM consists of $\mathcal{T}$ Transformer layers and a value function $\nu$. Perplexity (PPL) is usually used to evaluate LLMs in LLM pruning [31; 32; 9]. The lower the value, the better the performance. Therefore, the value function is denoted as $\nu(\mathcal{T})=1/\text{PPL}(\mathcal{T})$. Particularly, if there is no Transformer layer, the value function must be 0, i.e., $\nu(\emptyset)=0$. Let $\mathcal{S}\subseteq\mathcal{T}$ denote a subset of $\mathcal{T}$. According to [33], the SV of any Transformer layer $t\in\mathcal{T}$ is:
@@
where $w_{\mathcal{S}}=T\cdot\binom{T-1}{\mathcal{S}}=\frac{T!}{|\mathcal{S}|!(T-|\mathcal{S}|-1)!}$ is a coefficient. $\nu(\mathcal{S}\cup\{t\})-\nu(\mathcal{S})$ is known as Transformer layer $t$’s marginal contribution. Eq. (1) calculaties the marginal contribution for every subset $\mathcal{S}$, which results in a combinatorial explosion as the number of Transformer layers increases. Specifically, for $T$ Transformer layers, the total number of possible subsets $\mathcal{S}$ is $2^{T}$. For instance, with 32 Transformer layers in LLaMa-7B, the number of possible subsets becomes $2962^{32}=4,294,967,$ 296. The exact computation of Shapley values requires evaluating every subset that excludes the given Transformer layer, meaning that $32\times 2^{32}$ evaluations are needed. Each additional feature doubles the number of subsets that must be considered, causing an exponential increase in the number of computations. Consequently, this problem is NP-hard.